# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v3

@@ 4 Methodology
This work studies the spiking graph domain adaptation problem and proposes a new approach DeSGDA. DeSGDA consists of three parts: Degree-aware personalized spiking representation utilizes different thresholds for different degrees, effectively addressing the inflexible architecture challenge; Adversarial distribution alignment uses the adversarial training on membrane potential against a domain discriminatory to align distribution between different domains, and Pseudo-label distillation further applies the pseudo-label to enhance model performance. We provide the theoretical guarantee of DeSGDA to ensure the effectiveness. The overview of DeSGDA is shown in Figure 1.
Problem Setup. Given a graph $G=(V,E,\mathbf{X})$ with the node set $V$, the edge set $E$, and the node attribute matrix $\mathbf{X}$. Denote $S$ as the binary input sampled from Bernoulli distribution with probability of $\mathbf{X}$. The labeled source domain is denoted as $\mathcal{D}^{s}=\{(G_{i}^{s},y_{i}^{s})\}_{i=1}^{N_{s}}$, where $y_{i}^{s}$ denotes the labels of $G_{i}^{s}$; the unlabeled target domain is $\mathcal{D}^{t}=\{G_{j}^{t}\}_{j=1}^{N_{t}}$, where $N^{s}$ and $N^{t}$ denote the number of source graphs and target graphs. Both domains share the same label space $\mathcal{Y}$ but have different distributions in the graph space. We aim to train a spiking graph model using labeled source graphs and unlabeled target graphs to achieve superior performance in the target domain.
4.1 Degree-aware Personalized Spiking Representation
In this part, we first study the disadvantages of directly applying SNNs to graphs and then propose the degree-aware personalized spiking representation. Existing SGNs Li et al. (2023); Yao et al. (2023a); Duan et al. (2024) usually employ a global threshold for membrane potential firing. However, the global threshold can lead to the inflexible architecture issue since nodes with higher degrees are more likely to trigger spikes than those with lower degrees. As shown in Eq. 1, nodes with higher degrees have more neighbors, and the aggregation operation allows for more significant feature accumulation, making it easier for these nodes to trigger spikes compared to those with fewer neighbors. To alleviate this issue, we propose the degree-aware thresholds and iteratively update their values.
@@ 4.1 Degree-aware Personalized Spiking Representation
Specifically, we first $set$ all the degrees of nodes in the source domain graphs, i.e., $D^{s}=set(D_{1}^{s}\cup\cdots\cup D_{N_{s}}^{s})$, where $D_{i}$ denotes the degree set of graph $G_{i}^{s}$, and $set(\cdot)$ operation is an unordered sequence of non-repeating elements. Considering that low-degree nodes are more challenging to trigger while high-degree nodes trigger more easily, we propose setting higher thresholds for high-degree nodes and lower for low-degree nodes, which is formulated as:
$s_{\tau}^{d_{i}^{s}}=\mathbb{H}(u_{\tau}$ $-V_{th}^{d_{i}^{s}}),\quad S^{d_{i}^{s}}=avg(s_{\tau}^{d_{i}^{s}}),$ (2)
$V_{th}^{d_{i}^{s}}$ $=(1-\alpha)V_{th}^{d_{i}^{s}}+\alpha S^{d_{i}^{s}},$
where $V_{th}^{d_{i}^{s}}$ is the threshold of degree $d_{i}^{s}\in D^{s}$, initially set to $V_{th}$, and $\alpha$ is a hyper-parameter. The $avg(\cdot)$ operation takes the average of spiking representation with degree $d_{i}^{s}$. Consequently, high-degree nodes tend to achieve high $S^{d_{i}^{s}}$, which leads to an iterative increase in the threshold corresponding to degree $d_{i}^{s}$ and conversely for lower-degree nodes. To further explore the background mechanism of the hypothesis, we have the following analysis.
Hypothesis 1
In graph spiking networks, nodes with low-degree are more challenging to trigger while high-degree nodes trigger more easily.
The details analysis are introduced in Appendix A. With different thresholds for different node degrees, we can obtain the personalized node spiking representation $\mathbf{s}_{v\in G_{i}^{s}}^{d_{j}^{s}}$. Then, we summarize all node representations with a readout function into the graph-level representation and output the prediction with a multi-layer perception (MLP) classifier:
$\mathbf{s}_{i}=$ $\operatorname{READOUT}\left(\left\{\mathbf{s}_{v}^{d_{j}^{s}}\right\}_{v\in G_{i}^{s}}\right),$ (3)
$\hat{y}_{i}^{s}=H(\mathbf{s}_{i}),$
where $\hat{y}_{i}$ is the predicted result and $H(\cdot)$ is the classifier. After that, the source classification loss is:
$\mathcal{L}_{S}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}{l}(y_{i}^{s},\hat{y}_{i}^{s}),$ (4)
where $l(\cdot)$ is the loss function and $y_{i}^{s}$ is the ground truth of the $i$-th graph $G_{i}^{s}$ in the source domain.
However, in the scenario of domain adaptation, two significant issues remain unresolved. The first issue is that degrees in the target domain may be unseen in the source, rendering the thresholds ineffective for these degrees. The second is that the thresholds in the target may differ from those in the source, simply applying the source domain’s thresholds could lead to performance degradation. To alleviate the first issue, we initialize the threshold $V_{th}^{d_{i}^{t}}$ with the same value, where $d_{i}^{t}\notin D^{s}$. Then, with the training process of adversarial alignment, we iteratively update the threshold for degree $d_{i}^{t}$ with Eq. 2. To address the second issue, we incorporate the pseudo-label distillation module in Section 4.3 to guide the update of source degree thresholds on the target domain.
4.2 Adversarial Distribution Alignment
@@
To eliminate the discrepancy between the source and target domains, we propose the adversarial distribution alignment module. Specifically, for each source graph $G_{i}^{s}$ and target graph $G_{i}^{t}$, we use the degree-aware personalized spiking GNNs-based encoder $F(\cdot)$ and semantic classifier $H(\cdot)$ to produce predicted labels. Then, a domain discriminator $Q(\cdot)$ is trained to distinguish features from the source and target domains. The encoder and classifier are adversarial trained to align the feature spaces of the source and target domains.
$\mathcal{L}_{AD}$ $=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$
$+\mathbb{E}_{G_{j}^{t}\in\mathcal{D}^{t}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right).$
However, the degree in the target domain may be unseen by the source. Thus, we further initialize the threshold with $V_{th}^{d_{j}^{t}}$ and $d_{j}^{t}\notin D^{s}$, which is formulated as:
$\mathcal{L}_{AD}$ $=\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ D_{j}^{t}\subset{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right)$
$+\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$
@@
$+\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ \exists d_{j}^{t}\notin{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}},V_{th}^{D^{t}}\right)\right),$
@@
where $D^{t}=\{d_{i}^{t}|d_{i}^{t}\in D^{t},d_{i}^{t}\notin D^{s}\}$. We iteratively update $V_{th}^{D^{t}}$ with Eq. 2 on each latency. Furthermore, we present an upper bound for adversarial distribution alignment.
@@
Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Theorem 1, the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$ and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality :
$\epsilon_{T}(h,$ $\hat{h}_{T}(\mathbf{X}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$
$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+\min\Big{(}|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,$
$|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\Big{)}+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right),$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h}(\mathbf{S}))=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(\mathbf{S}_{n})-\hat{h}(\mathbf{S}_{n})|$ and $\epsilon_{T}(h,\hat{h}(\mathbf{X}))=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h}(\mathbf{X}))+\epsilon_{T}(h,\hat{h}(\mathbf{X}))\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
$\mathbf{P}_{kt}=\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}$ (5)
Theorem 2 proves the generalization bound of spiking graph domain adaptation. More details can be found in Appendix B.
@@ 4.3 Pseudo-label Distillation for Discrimination Learning
To further address the variance in thresholds between the target and source domains, we incorporate the pseudo-label distillation module into the DeSGDA framework. With reliable pseudo-labels, we can effectively update the source degree thresholds in the target domain.
The goal of the pseudo-label distilling procedure is to keep those examples and their corresponding pseudo-labels from the deep feature space that aligns with the shallow feature space. Specifically, we denote $\mathbf{s^{\prime}}_{i}^{t}$ as the shallow spiking graph representation on the $L^{\prime}$-th layer, where $L^{\prime}<L$, and $\hat{y}_{i}^{t}$ as the prediction of graph $G_{i}^{t}$ on the $L$-th layer. Then, to enhance alignment between the shallow and deep feature spaces and facilitate the generation of more accurate pseudo-labels, we cluster the shallow features $\mathbf{s^{\prime}}^{t}$ into $C$ clusters and each cluster $\mathcal{E}_{j}$ includes graphs $\{G_{j}^{t}\}$. After that, we find the dominating labels $e_{r}$ in the cluster, i.e., $\max_{r}|\{\mathcal{E}_{r}:e_{r}=\hat{y}_{j}^{t}\}|$, and remove other instances with the same pseudo-label but in different clusters. Formally, the pseudo-labels are signed with:
$\mathcal{P}=\left\{\left({G}_{j}^{t},\hat{y}_{j}^{t}\right):e_{j}=\max_{r}\left|\left\{\mathcal{E}_{r}:e_{r}=\hat{y}_{j}^{t}\right\}\right|\right\}.$ (6)
@@
$\mathcal{L}_{T}=\mathbb{E}_{{G}_{j}^{t}\in\mathcal{P}}l\left({H(\mathbf{s}}_{j}^{t}),\hat{y}_{j}^{t}\right),$ (7)
where $H(\cdot)$ and $\mathbf{s}_{j}^{t}$ are the classifier and spiking graph representation, which are defined in Eq. 3. $l(\cdot)$ is the loss function, and we implement it with cross-entropy loss.
Theorem 3
Under the assumption of Theorem 1, we further assume that there exists a small amount of i.i.d. samples with pseudo labels $\{(G_{n},Y_{n})\}_{n=1}^{N^{\prime}_{T}}$ from the target distribution $\mathbb{P}_{T}(G,Y)$ $(N^{\prime}_{T}\ll N_{S})$ and bring in the conditional shift assumption that domains have different labeling function $\hat{h}_{S}\neq\hat{h}_{T}$ and $\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})}=C_{h}\leq C_{f}C_{g}(D\in\{S,T\})$ for some constant $C_{h}$ and distance measure $\eta$, and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i}^{s},y_{i}^{s})\}_{i\in[n]}$, with probability at least $1-\delta$ the following inequality holds:
$\epsilon_{T}(h,\hat{h}_{T}($ $\mathbf{X}))\leq\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\cdot\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right)$
$+\min\Big{(}|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,$
$|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\Big{)}\Bigg{)}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$
$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right)+\omega^{\prime},$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Theorem 2.
The proof is detailed in Appendix C. From Theorem 3, we observe that the bound of DeSGDA is lower than simply aligning the distributions by incorporating the highly reliable pseudo-labels, demonstrating the effectiveness of pseudo labels for spiking graph domain adaptation.

### v4

@@ 4 Proposed Methodology
This paper proposes a novel framework DeSGraDA for the problem of spiking graph domain adaptation. DeSGraDA consists of three parts: Degree-Conscious Spiking Representation (Section 4.1), which assigns adaptive firing thresholds based on node degrees to address the limitations of rigid, fixed-threshold architectures; Temporal Distribution Alignment (Section 4.2), which employs adversarial training on temporal membrane potentials against a domain discriminator to align spiking dynamics across domains; and Pseudo-label Distillation further applies the pseudo-label to enhance model performance. We also provide a theoretical generalization bound to support the effectiveness of DeSGraDA. An overview of the framework is in Figure 1.
4.1 Degree-Conscious Spiking Representation
First, we study the disadvantages of directly applying SNNs to graphs and then propose the degree-conscious spiking representation module. Existing SGNs (Li et al., 2023; Duan et al., 2024) usually employ a fixed global threshold ($V_{th}$ in Eq. 1 and 2) for firing. Assume that the membrane potential of node $u_{\tau,i}$ follows the normal distribution $\mathcal{N}(\mu,\sigma^{2})$ (Kipf and Welling, 2016).
The following Proposition shows that the high-degree nodes are more likely to trigger spikes than the low-degree ones. Proof is in Appendix C.
Proposition 1
With aggregation operation in SGNs (i.e., $\mathcal{A}$ in Eq. 1), the expectation of the updated node membrane potential is: $\mathbb{E}(u_{\tau+1,i})\sim\mathcal{N}\left(\left(1+\sum\nolimits_{j\in N(i)}w_{ij}\right)\mu,\left(1+\sum\nolimits_{j\in N(i)}w_{ij}\right)^{2}\sigma^{2}\right)$, where $N(i)$ is the set of node $i$’s neighbors, and $w_{ij}$ is the weight between nodes $i$ and $j$.
Figure 2: Correlation between spiking frequency and aggregation weights under fixed thresholds on the PROTEINS and DD datasets.
From Proposition 1, we observe that node $i$ follows a normal distribution with a mean of $(1+\sum_{j\in N(i)}w_{ij})\mu$, determined by the aggregated weights of its neighboring nodes [^2] This does not account for models that constrain the sum of the neighboring weights (e.g., GAT (Veličković et al., 2018)).. We conduct an experiment to investigate the relationship between the aggregated weight ($\sum\nolimits_{j\in N(i)}w_{ij}$) and spiking frequency (i.e., count of $u_{T}>V_{th}$) for a fixed threshold $V_{th}$. As shown in Figure 2, the spiking frequency and aggregation weights under fixed thresholds exhibit a relatively high correlation coefficient, indicating that nodes with higher degrees tend to accumulate more features from neighbors, making it easier to trigger spikes than those with fewer neighbors. Fixed thresholds inherently bias spiking activation toward high-degree nodes, leading to under-representation of structurally important yet sparsely connected nodes and undermining both expressiveness and generalization in SGNs.
@@ 4.1 Degree-Conscious Spiking Representation
To alleviate this problem, we propose the use of degree-conscious thresholds. Specifically, let $B_{i}^{s}$ be the set of degrees for the nodes in graph $G_{i}^{s}$. We collect all unique node degrees across the source domain graphs with $B^{s}=set(B_{1}^{s}\cup\cdots\cup B_{N_{s}}^{s})$, where $set(\cdot)$ operation is an unordered sequence of distinct elements. With the observation from Proposition 1, we propose setting higher thresholds for high-degree nodes and lower for low-degree nodes. For node $v$ with degree $d_{v}^{s}$, we have:
$\;\;s_{\tau,v}^{d_{v}^{s}}=\mathbb{H}(u_{\tau,v}-V_{th}^{d_{v}^{s}}),\quad V_{th,\tau+1}^{d_{v}^{s}}=(1-\alpha)V_{th,\tau}^{d_{v}^{s}}+\alpha\bar{s}^{d_{v}^{s}}_{\tau,v},$ (3)
$u_{\tau+1,v}=\lambda(u_{\tau,v}-V_{th}s_{\tau,v}^{d_{v}^{s}})+\sum\nolimits_{j\in N(v)}w_{ij}\mathcal{A}\left(A,s_{\tau,v}^{d_{j}^{s}}\right)+b_{i},$
where $V_{th}^{d_{v}^{s}}$ is the threshold for nodes with degree $d_{v}^{s}\in B^{s}$, initially set to $V_{th}$, and $\alpha$ is a hyper-parameter. $\bar{s}^{d_{v}^{s}}_{\tau}$ is the average of spiking representation $s_{\tau}^{d_{v}^{s}}$ with degree $d_{v}^{s}$ on latency step $\tau$. In Eq. 3, we dynamically update the threshold $V^{d^{s}_{v}}_{th,\tau+1}$ with $(1-\alpha)V^{d^{s}_{v}}_{th,\tau}+\alpha\bar{s}^{d^{s}_{v}}_{\tau,v}$ based on the average spiking frequency $\bar{s}^{d^{s}_{v}}_{\tau,v}$ of nodes with degree $d^{s}_{v}$. Consequently, high-degree nodes typically exhibit higher $\bar{s}^{d^{s}_{v}}_{\tau,v}$, resulting in increased thresholds $V^{d^{s}_{v}}_{th,\tau+1}$ to suppress over-activation, while low-degree nodes yield lower thresholds to enhance activation, enabling adaptive spiking control across structurally diverse nodes. For each node $v$ in graph $G_{i}^{s}$, we calculate the membrane potential $u_{\tau,v}^{s}$, and summarize all the node representations with a readout function (Xu et al., 2018; Wang et al., 2024a) into the graph-level representation:
$\mathbf{s}_{G_{i}^{s}}=\operatorname{READOUT}\left(\left\{{s}_{T}^{d_{v}^{s}}\right\}_{v\in G_{i}^{s}}\right),$ (4)
where $T$ is the total number of latency steps. Finally, we output the prediction $\hat{y}_{G_{i}^{s}}^{s}=H(\mathbf{s}_{G_{i}^{s}})$ with a classifier $H(\cdot)$ by minimizing the source classification loss $\mathcal{L}_{S}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}{l}(y_{G_{i}^{s}}^{s},\hat{y}_{G_{i}^{s}}^{s})$, where $l(\cdot)$ is the loss function and $y_{i}^{s}$ is the ground truth of graph $G_{i}^{s}$ in the source domain.
4.2 Temporal Distribution Alignment
Unlike GNNs, spiking models rely on membrane potential dynamics to generate sparse spike trains, making their spike representations highly dynamic and non-differentiable. Existing GDA (Yin et al., 2023; Chen et al., 2025) methods assume continuous, static embeddings and thus fail to align the time-dependent neural dynamics in SGNs. To address this, we propose a temporal-based alignment framework that captures and matches the evolution of spiking patterns across domains.
First, we introduce a temporal attention mechanism that adaptively aggregates time-dependent neuronal states. Specifically, given the temporal membrane potential of graphs on each latency step, we have $[\mathbf{u}_{1,G_{i}},\ldots,\mathbf{u}_{\tau,G_{i}},\cdots,\mathbf{u}_{T,G_{i}}]$, where $\mathbf{u}_{\tau,G_{i}}=\operatorname{READOUT}(\left\{{u}_{\tau,v}\right\}_{v\in G_{i}})$ and $G_{i}\in\{\mathcal{D}^{s},\mathcal{D}^{t}\}$. We stack the $T$ steps membrane potential into $\mathbf{U}_{G_{i}}\in\mathbb{R}^{T\times d}$, where $d$ is the hidden dimension. The goal is to learn an importance-weighted $\alpha_{\tau}$ to summarize of temporal membrane potential representation, which is formulated as follows:
$\alpha_{\tau}=\operatorname{Atten}(\mathbf{U}_{G_{i}}),\quad\tilde{\mathbf{U}}_{G_{i}}=\sum\nolimits_{\tau=1}^{T}\alpha_{\tau}\mathbf{u}_{\tau,G_{i}},$ (5)
@@ 4.2 Temporal Distribution Alignment
where $\operatorname{Atten}(\cdot)$ is the self-attention operation (Vaswani et al., 2017). Then, we propose the adversarial distribution alignment module to eliminate the discrepancy between the source and target domains. Specifically, for each source graph $G_{i}^{s}$ and target graph $G_{j}^{t}$, we denote the semantic classifier as $H(\cdot)$ to produce predicted labels, and a domain discriminator $Q(\cdot)$ to distinguish features from the source and target domains. The temporal-based distribution alignment module is adversarially trained to align the feature spaces of the source and target domains, which is formulated as:
@@
$\mathcal{L}_{AD}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(H(\tilde{\mathbf{U}}_{G_{i}}^{s})|V_{th}^{B^{s}}\right)+\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t},\exists d_{j}^{t}\notin{B^{s}}\end{subarray}}\log\left(1-Q\left(H(\tilde{\mathbf{U}}_{G_{i}}^{t})|V_{th}^{B^{s}},V_{th}^{B^{t}}\right)\right),$
@@
where $B^{t}=\{d_{i}^{t}|d_{i}^{t}\in B^{t},d_{i}^{t}\notin B^{s}\}$. We iteratively update $V_{th}^{B^{t}}$ with Eq. 3 on each latency. Furthermore, we present an upper bound for temporal-based distribution alignment.
@@
Assume that the learned discriminator is $C_{g}$-Lipschitz continuous, the feature extractor $f$ is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$ and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the inequality :
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (6)
$+\omega+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
[…]
The proof is proposed in Appendix D. Theorem 2 justifies that the generalization gap of spiking GDA relies on the domain divergence $2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))$ and model discriminability $\omega$, as well as the model’s ability to avoid overfitting to the training data, which is quantified by the Rademacher complexity term $2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$. In the application of spiking GDA, this term captures how the model’s sensitivity to random fluctuations in the node feature aggregation (especially for higher-degree nodes) can lead to overfitting, thus affecting the model’s ability to generalize to the target domain. This overfitting risk is particularly relevant when the model is too flexible in fitting the training data, exacerbating the generalization gap, especially under domain shifts.
@@ 4.3 Pseudo-label Distillation for Discrimination Learning
To further tiny the generalization gap between the target and source domains, we incorporate the pseudo-label distillation module into the DeSGraDA framework. The goal of the module is to ensure consistent prediction between the shallow and deep layers. Specifically, let $\mathbf{s^{\prime}}_{\tau,G_{i}}^{t}$ be the shallow spiking graph representation of $G_{i}$ on the latency step $\tau$ ($\tau<T$) in the target domain, and $\hat{y}_{G_{i}}^{t}=H(\mathbf{s}_{G_{i}})$ be the prediction of graph $G_{i}$. Then, to enhance consistency between the shallow and deep feature spaces and facilitate the generation of more accurate predictions, we cluster the shallow graphs features in the target domain into $C$ clusters, and each cluster $\mathcal{E}_{j}$ includes graphs $\{G_{j}^{t}\}$. After that, we find the dominating labels $e_{r}$ in the cluster, i.e., $\max_{r}|\{\mathcal{E}_{r}:e_{r}=\hat{y}_{G_{j}}^{t}\}|$, and remove other samples with the same prediction but in different clusters. Formally, the pseudo-labels are signed as:
$\mathcal{P}=\left\{\left({G}_{j}^{t},\hat{y}_{G_{j}}^{t}\right):e_{j}=\max_{r}\left|\left\{\mathcal{E}_{r}:e_{r}=\hat{y}_{G_{j}}^{t}\right\}\right|\right\}.$ (7)
@@
$\mathcal{L}_{T}=\mathbb{E}_{{G}_{j}^{t}\in\mathcal{P}}l\left({H(\mathbf{s}}_{G_{j}^{t}}^{t}),\hat{y}_{G_{j}^{t}}^{t}\right),$ (8)
where $\mathbf{s}_{G_{j}^{t}}^{t}$ is the spiking graph representation of $G_{j}^{t}$ in the target domain. $l(\cdot)$ is the loss function, and we implement it with cross-entropy loss. We further analyze the generalization bound by applying the pseudo-label distillation module, and the proof is detailed in Appendix E. From the proof, we observe that the bound is lower than simply aligning the distributions by incorporating the highly reliable pseudo-labels, demonstrating the effectiveness of pseudo-labels for spiking graph domain adaptation.